# citations.bib

### v2

@@ III-D Evaluation Metrics
To comprehensively assess MARCO’s performance and utility, we adopted a multi-dimensional evaluation framework, spanning execution performance, code quality, and cost-efficiency:
  • Performance: We measured execution time, floating-point operations per second (FLOPs), and memory usage. Specifically: Execution time was measured using system-level profiling tools, such as time.h in C and high-resolution timers available in Python, following best practices established in LLM benchmarking literature [li2024performance]. We used Floating Point Operations Per Second (FLOPS) to perform direct comparisons of computational throughput across different optimization strategies. Memory usage was assessed using packages like tracemalloc [tracemalloc] (for tracking high-memory objects) and psutil [psutil] (for monitoring system-level resource utilization), consistent with techniques reported in recent HPC optimization studies [ding2023hpc, yin2023ion].
  • Code Quality: We evaluated correctness and readability. Correctness was validated by submitting generated solutions to the LeetCode problem submission framework, which provides automated feedback on pass/fail test cases. While certain limitations existed due to library import restrictions, this approach aligned with established methods in prior code-generation evaluations [li2024performance]. For user-provided code, MARCO aimed to preserve original structure and readability, making localized improvements rather than complete rewrites. All modifications were accompanied by inline comments and high-level explanations to maintain developer transparency.
  • Cost: We measured API cost per query and token consumption: m During March and April, the Anthropic (Claude) models processed approximately 562,703 input tokens and 213,227 output tokens, incurring a cost of \$4.22 as recorded in the Anthropic developer console. For OpenAI (ChatGPT) models over February to April, 1,145,708 input tokens and 898,465 output tokens were consumed, at a total cost of \$10.50, yielding an average API cost per query of approximately \$0.00369. These cost assessments are critical for evaluating the practicality of multi-agent frameworks like MARCO, especially relative to large-scale fine-tuning or dedicated model retraining [nichols2023hpc].
@@ V Experimental Setup
To ensure reproducibility and provide a fair evaluation of MARCO’s performance, all experiments were conducted under controlled hardware and software conditions. Specifically, we ran benchmarks on a machine equipped with an AMD EPYC 7502 32-Core Processor, NVIDIA A100 GPUs, and 256 GB of system memory, operating under Ubuntu 22.04 LTS. The software environment included Python 3.10, CUDA 12.2 (for GPU-accelerated tasks), and relevant package dependencies such as tracemalloc version 3.11 and psutil version 5.9.8.
@@
Each experiment was repeated N = 5 times to account for random variability, and reported metrics (e.g., execution time, memory usage) reflect the average over these runs, with standard deviations provided where applicable. Random seeds were fixed at 42 where supported to maintain consistency. All code was executed within a venv virtual environment to ensure consistent dependency resolution across experiments. Additionally, Docker containers (python:3.10-slim with CUDA 12.2 support) were used in select experiments to support dynamic library installation and compatibility, particularly when testing code requiring packages beyond the default system installation.
V-A Performance Analysis
@@
Figure 1: Runtime distribution comparison between problems of Easy and Medium difficulty levels from the LeetCode 75 dataset and 10 selected problems of Hard difficulty.
@@
We systematically evaluated MARCO-optimized code against baseline LLM-generated code (including Claude-3.5 Sonnet, GPT-4o, DeepSeek Coder-V2, and Llama 3.1 8b) on a diverse set of programming tasks from the LeetCode 75 problem set, as well as 10 additional randomly selected hard problems. Our analysis focused on runtime performance, consistency, and relative improvements. We report results from comparisons with Claude-3.5 Sonnet, which consistently exhibited higher code performance than the other tested LLMs.
@@
Table I summarizes average runtime across models.
TABLE I: Average Runtime (ms) Across Models and Problem Difficulties
Model Easy Medium Hard
Claude-3.5 Sonnet $11.811.8$ $78.178.1$ $288.1288.1$
MARCO (ours) $9.39.3$ $65.965.9$ $138.5138.5$
V-B Memory Usage Analysis
We compared memory usage between MARCO-optimized and baseline LLM-generated code using peak memory consumption metrics (recorded using tracemalloc, psutil, or [MEMORY_TOOL]). Among the 85 evaluated problems (75 easy/medium, 10 hard), MARCO showed noticeable memory usage improvements in $1515$ cases, while the baseline LLMs showed advantages in $1010$ cases, with the remaining $6060$ showing comparable performance.
[…]
Table II presents memory usage comparisons.
TABLE II: Average Peak Memory Usage (MB) Across Models
Model Easy Medium Hard
Claude-3.5 Sonnet $18.918.9$ $25.325.3$ $32.332.3$
MARCO (ours) $18.918.9$ $25.225.2$ $28.828.8$
V-C Summary
Overall, MARCO demonstrated consistent runtime improvements on hard problems, competitive performance on medium problems, and robust memory efficiency across a diverse range of tasks. These findings establish MARCO as a promising framework for scalable, architecture-aware code optimization in HPC workloads, warranting further exploration into cross-model comparisons, reinforcement-based optimization loops, and domain-specific extensions.
VI Multi-Agent System for Optimizing LLM Code (MARCO)
VI-A Motivation
The rapid growth of computational demands across industries has driven unprecedented interest in high-performance computing (HPC) applications. As the large language model (LLM) era unfolds, there has been a parallel surge in attempts to leverage LLMs for HPC-related code generation and optimization. However, our findings reveal that state-of-the-art LLMs exhibit notable inefficiencies when generating HPC code, motivating the design of a multi-agent system (MAS).
A key limitation arises from the pretraining knowledge cut-off of current LLMs, which restricts their access to the latest advances in HPC architectures, optimization strategies, and software libraries. To address this gap, we integrate a Web-Search component into our agent system, enabling dynamic retrieval of up-to-date optimization techniques from scholarly resources such as IEEE, ACM, arXiv, and ResearchGate. This allows MARCO to explore and incorporate emerging methods, language updates, and cutting-edge toolchains, surpassing the static knowledge of the underlying LLM.
Another core motivation is cost-efficiency. Prior work on fine-tuning monolithic LLMs for HPC code generation demands extensive GPU resources and substantial training overhead [ding2023hpc]. Manual tuning, while often effective, requires deep expert knowledge and is labor-intensive, raising the barrier to entry for scientific software developers. MARCO leverages pretrained LLMs without the need for fine-tuning, instead enriching them with system-level context and externally sourced insights, making the solution both scalable and accessible. By combining up-to-date research knowledge with automated code transformation and iterative feedback, MARCO aims to significantly reduce the need for manual optimization and deliver diverse, performant solutions for large-scale HPC projects, ultimately lowering the barrier to innovation in high-performance software engineering.
VI-B MARCO Framework
Figure 2: MARCO Pipeline Overview. Interaction between code optimizer, web search engine, and performance evaluator agents in the iterative optimization loop.
The MARCO framework is composed of three core components, working in a coordinated loop:
  • Web-Search Engine: We integrate the Tavily search tool [tavily2024] to retrieve relevant techniques from scholarly sources, including acm.org, ieee.org, arxiv.org, and researchgate.net. The search engine is configured with adjustable depth and result count parameters (search_depth, include_answer, N sources per iteration) - where N is the number of user-selected iterations, ensuring that the optimizer agent has access to diverse and up-to-date references for each optimization pass.
    Figure 3: Web-Search Integration. Schematic showing how external scholarly sources feed optimization strategies into MARCO’s loop.
  • Performance Evaluator Agent: This agent benchmarks the generated code across multiple metrics, including execution time, memory usage, algorithmic complexity, and computational cost. It provides structured feedback to the optimizer agent, allowing the system to iteratively refine solutions. Performance data is presented to users in a tabular front-end, summarizing performance trends across iterations, correctness evaluations, and complexity assessments.
VI-C Evaluation of MARCO
We evaluated MARCO-optimized code against baseline LLM-generated solutions using the LeetCode 75 dataset [leetcode75] and an additional set of 10 randomly selected hard problems. We specifically compared the performance of MARCO with and without the Web-Search component, benchmarking results against Claude-3.5 Sonnet and ChatGPT (GPT-4o).
Across all tests, MARCO with the Tavily Web-Search integration consistently outperformed both Claude-3.5 Sonnet and MARCO without the web-search enhancement. This reflects the growing importance of incorporating real-time, externally sourced optimization strategies, especially in the fast-evolving HPC landscape, where new architectures, libraries, and techniques frequently emerge.
We observed that iterative refinement contributed significantly to performance gains: on harder LeetCode problems, the final iteration often achieved the best results, showing an average improvement of 30.9% over earlier iterations. Figure 4 illustrates the runtime differences between MARCO with and without the Tavily web-search tool across the evaluated problem set.
Figure 4: Runtime Comparison. Runtime differences between MARCO with and without the Tavily Web-Search tool on the LeetCode 75 problem set.
VII Discussion & Insights
VII-A Key Observations
Our experiments reveal that MARCO substantially enhances execution speed, memory efficiency, and parallelism compared to baseline LLM-generated code. Notably, the integration of the Web-Search tool consistently yields superior optimizations, enabling the system to overcome the inherent knowledge limitations of the underlying LLMs. By dynamically retrieving cutting-edge techniques and research, MARCO bridges the gap between static pretrained models and the evolving landscape of HPC optimization.
VII-B Limitations
Despite its advantages, MARCO introduces trade-offs between multi-agent query complexity and inference latency. Specifically, the iterative communication between agents and the need to transmit intermediate outputs result in slower code generation compared to traditional single-pass LLM workflows. Additionally, although the Web-Search component enables access to recent research, the system’s performance is still constrained by the LLM’s ability to correctly interpret and implement advanced HPC techniques.
At present, the Performance Evaluator Agent lacks native support for several key HPC libraries, including Numba, Torch, and CUDA, limiting the system’s optimization potential for GPU-accelerated workloads. Furthermore, MARCO’s current implementation is restricted to Python due to its rich ecosystem of optimization libraries and tooling. Another noteworthy limitation is cost variability: while ChatGPT (GPT-4o) offers advanced reasoning capabilities, it incurs higher API costs and, in our benchmarks, occasionally underperformed relative to Claude 3.5 Sonnet. These findings suggest opportunities for future work to enhance agent reasoning, expand language and library support, and optimize cost-performance trade-offs.
VII-C Unexpected Findings
An intriguing observation emerged during our benchmarking on the LeetCode 75 dataset: in certain cases, the standalone LLM outperformed MARCO, particularly on ‘easy’ problems. Specifically, the LLM achieved better runtime and memory performance on trivial tasks, where MARCO’s multi-agent orchestration introduced unnecessary computational overhead. One hypothesis is that some of these problems may have been included in Claude 3.5’s training data, providing it with a prior advantage. However, since MARCO also leverages Claude 3.5 as part of its pipeline, this alone does not fully explain the discrepancy, especially given the lack of transparency around model training datasets. A more compelling explanation is that simple problems inherently require minimal optimization, and thus, MARCO’s iterative refinement process introduces superfluous complexity. Consistent with our system design goals, our results show that as problem difficulty increases, MARCO’s advantage over the standalone LLM becomes increasingly pronounced, highlighting its strength in accelerating complex and computationally demanding tasks.
VIII Conclusion & Future Work
In this work, we benchmarked the performance of LLM-generated HPC code and introduced MARCO, a novel multi-agent system designed to iteratively refine code for improved performance, memory efficiency, and scalability. Our results demonstrate that MARCO effectively leverages external knowledge sources and iterative optimization loops to outperform standalone LLMs, particularly on complex and computationally demanding tasks.
Despite these promising results, substantial opportunities remain for further development and testing. Enhancing both system capabilities and evaluation rigor is essential to ensure that MARCO consistently delivers state-of-the-art performance. One important direction involves improving the communication between the Code Optimizer Agent and the Performance Evaluator Agent. Currently, the testing agent provides feedback solely in the form of final performance metrics; however, richer, bidirectional communication, such as intermediate diagnostic signals or targeted optimization hints, could significantly improve the refinement process. We also envision integrating Docker-based execution environments to support a wider range of external libraries (e.g., Numba, Torch, CUDA) and enable more robust, cross-platform testing.
Another key improvement area is the optimization workflow itself. Rather than generating a new code variant for each iteration, MARCO could benefit from maintaining and editing a persistent artifact, enabling finer-grained incremental improvements and reducing redundant computations. Additionally, extending MARCO’s capabilities to support advanced cost metrics and smarter, context-aware improvement recommendations could further enhance its optimization effectiveness.
Incorporating reinforcement learning techniques, as demonstrated in systems like AlphaCode [li2022competition], could further enhance MARCO’s optimization capabilities by enabling it to learn from iterative feedback and improve over time. Integrating reinforcement learning to guide the agent’s search and decision-making processes [sun2023reinforcement, sun2024llm] presents another promising avenue, potentially improving both convergence speed and solution quality. Furthermore, enhancing MARCO with persistent communication history across sessions—akin to LLM chat memory—could reduce redundant token usage, improve operational cost-efficiency, and enable richer multi-session optimization pipelines.
Overall, MARCO represents an important step toward automated, scalable HPC code optimization, with the potential to significantly lower the barrier to entry for high-performance programming and accelerate innovation across computational disciplines.
In summary, MARCO advances the frontier of automated HPC code optimization by combining multi-agent orchestration, real-time knowledge retrieval, and iterative refinement. By bridging the gap between static pretrained LLMs and the rapidly evolving landscape of high-performance computing, MARCO sets the stage for scalable, cost-efficient, and accessible optimization pipelines that can accelerate discovery and innovation across scientific and industrial domains.
[…]

### v3

@@ III-D Evaluation Metrics
To comprehensively assess MARCO’s performance and utility, we adopted a multi-dimensional evaluation framework spanning execution performance, code quality, and cost-efficiency:
  • Performance Metrics: We measured three key performance indicators:
    • Execution time was measured using system-level profiling tools, such as time.h in C and high-resolution timers in Python, following best practices established in LLM benchmarking literature [li2024performance].
    • Computational throughput was assessed using FLOPS measurements to enable direct comparisons across different optimization strategies.
    • Memory usage was evaluated using Python packages including tracemalloc for tracking high-memory objects and psutil for monitoring system-level resource utilization, consistent with techniques reported in recent HPC optimization studies [ding2023hpc, yin2023ion].
  • Code Quality Assessment: We evaluated two primary aspects:
    • Correctness was validated by submitting generated solutions to the LeetCode problem framework, which provides automated feedback on test case performance. While certain limitations existed due to library import restrictions, this approach aligned with established methods in prior code-generation evaluations [li2024performance].
    • Readability and maintainability were preserved through MARCO’s approach of making localized improvements rather than complete rewrites. All modifications were accompanied by inline comments and high-level explanations to maintain developer transparency.
  • Cost Analysis: We measured API cost per query and token consumption. During March and April, the Anthropic (Claude) models processed approximately 562,703 input tokens and 213,227 output tokens, incurring a cost of \$4.22 as recorded in the Anthropic developer console. For OpenAI (ChatGPT) models over February to April, 1,145,708 input tokens and 898,465 output tokens were consumed, at a total cost of \$10.50, yielding an average API cost per query of approximately \$0.00369. These cost assessments are critical for evaluating the practicality of multi-agent frameworks like MARCO, especially relative to large-scale fine-tuning or dedicated model retraining [nichols2023hpc].
@@ IV-A Experimental Setup
To ensure reproducibility and provide a fair evaluation of MARCO’s performance, all experiments were conducted under controlled hardware and software conditions. Specifically, we ran benchmarks on a machine equipped with an AMD EPYC 7502 32-Core Processor, NVIDIA A100 GPUs, and 256 GB of system memory, operating under Ubuntu 22.04 LTS. The software environment included Python 3.10, CUDA 12.2 for GPU-accelerated tasks, and relevant dependencies including Python’s built-in tracemalloc module and psutil version 5.9.8.
@@
Each experiment was repeated 5 times to account for random variability, and reported metrics (execution time, memory usage) reflect the average over these runs, with standard deviations provided where applicable. Random seeds were set to 42 for all stochastic operations to maintain consistency. All code was executed within a venv virtual environment to ensure consistent dependency resolution across experiments. Additionally, Docker containers (python:3.10-slim with CUDA 12.2 support) were used in select experiments to support dynamic library installation and compatibility, particularly when testing code requiring packages beyond the default system installation.
IV-B Performance Analysis
@@
Figure 1: Runtime distribution comparison between Easy/Medium difficulty problems from the LeetCode 75 dataset and 10 selected Hard difficulty problems, showing MARCO’s performance improvements across difficulty levels.
@@
We systematically evaluated MARCO-optimized code against baseline LLM-generated code, including Claude-3.5 Sonnet, GPT-4o, DeepSeek Coder-V2, and Llama 3.1 8b, on a diverse set of programming tasks from the LeetCode 75 problem set and 10 additional randomly selected hard problems. Our analysis focused on runtime performance, consistency, and relative improvements. We report detailed results from comparisons with Claude-3.5 Sonnet, which consistently exhibited the highest code performance among the tested LLMs.
On the LeetCode 75 (easy and medium) problems, MARCO matched or outperformed Claude-3.5 Sonnet in 44 out of 75 cases, achieving an average runtime reduction of 14.6% (mean runtime 49.1 ms vs. 57.5 ms). For hard problems, MARCO demonstrated stronger gains, outperforming Claude-3.5 Sonnet in all 10 cases, with an average runtime improvement of 51.9%. Notably, MARCO’s solutions ranked in the top 20% of all LeetCode submissions for 7 out of 10 hard problems, demonstrating competitive performance against the broader programming community.
Table I summarizes average runtime performance across problem difficulties for the two best-performing approaches.
TABLE I: Average Runtime (ms) Across Problem Difficulties
Model Easy Medium Hard
Claude-3.5 Sonnet 11.8 78.1 288.1
MARCO (ours) 9.3 65.9 138.5
Improvement 21.2% 15.6% 51.9%
The results demonstrate that MARCO’s multi-agent approach becomes increasingly effective as problem complexity increases, with the most significant improvements observed on hard problems where sophisticated optimization strategies provide the greatest benefit.
IV-C Memory Usage Analysis
We compared memory usage between MARCO-optimized and baseline LLM-generated code using peak memory consumption metrics recorded with Python’s tracemalloc module and psutil. Among the 85 evaluated problems (75 easy/medium, 10 hard), MARCO showed measurable memory usage improvements in 15 cases, while baseline LLMs showed advantages in 10 cases, with the remaining 60 cases showing comparable performance within measurement variance.
On average, MARCO reduced peak memory consumption by 1.2%, with maximum observed savings of 8 MB on hard problems. Importantly, in some cases, MARCO’s optimizations slightly increased memory usage (by up to 1.9%) to achieve runtime speedups, highlighting the deliberate trade-offs between time and space optimizations that characterize effective HPC code.
Table II presents memory usage comparisons across problem difficulties.
TABLE II: Average Peak Memory Usage (MB) Across Problem Difficulties
Model Easy Medium Hard
Claude-3.5 Sonnet 18.9 25.3 32.3
MARCO (ours) 18.9 25.2 28.8
Change 0.0% -0.4% -10.8%
The memory analysis reveals that MARCO’s primary strength lies in runtime optimization rather than memory reduction, with the most significant memory improvements occurring on hard problems where algorithmic optimizations can substantially reduce space complexity.
IV-D Summary
Overall, MARCO demonstrated consistent and substantial runtime improvements on hard problems (51.9% average improvement), competitive performance on medium problems (15.6% improvement), and modest gains on easy problems (21.2% improvement). While memory optimization showed less dramatic improvements, MARCO maintained comparable or better memory efficiency across all problem categories. These findings establish MARCO as a promising framework for scalable, architecture-aware code optimization in HPC workloads, with particular effectiveness on complex algorithmic challenges. The results warrant further exploration into cross-model comparisons, reinforcement-based optimization loops, and domain-specific extensions to fully realize the potential of multi-agent code optimization approaches.
V Multi-Agent System for Optimizing LLM Code (MARCO)
V-A Motivation
The rapid growth of computational demands across industries has driven unprecedented interest in high-performance computing (HPC) applications. With the emergence of large language models (LLMs), there has been a parallel surge in attempts to leverage LLMs for HPC-related code generation and optimization. However, our findings reveal that state-of-the-art LLMs exhibit notable inefficiencies when generating HPC code, motivating the design of a multi-agent system (MAS).
Three key limitations drive our approach. First, a significant constraint arises from the pretraining knowledge cutoff of current LLMs, which restricts their access to the latest advances in HPC architectures, optimization strategies, and software libraries. To address this gap, we integrate a web-search component into our agent system, enabling dynamic retrieval of up-to-date optimization techniques from scholarly resources such as IEEE, ACM, arXiv, and ResearchGate. This allows MARCO to explore and incorporate emerging methods, language updates, and cutting-edge toolchains, surpassing the static knowledge of the underlying LLM.
Second, cost-efficiency presents a major concern. Prior work on fine-tuning monolithic LLMs for HPC code generation demands extensive GPU resources and substantial training overhead [ding2023hpc]. MARCO leverages pretrained LLMs without requiring fine-tuning, instead enriching them with system-level context and externally sourced insights, making the solution both scalable and accessible.
Third, accessibility barriers limit HPC adoption. Manual optimization, while often effective, requires deep expert knowledge and is labor-intensive, raising the barrier to entry for scientific software developers. By combining up-to-date research knowledge with automated code transformation and iterative feedback, MARCO aims to significantly reduce the need for manual optimization and deliver diverse, performant solutions for large-scale HPC projects, ultimately lowering the barrier to innovation in high-performance software engineering.
V-B MARCO Framework
Figure 2: MARCO Pipeline Overview. The iterative optimization process showing the interaction between the code optimizer agent, web search engine, and performance evaluator agent. The system begins with input code, searches for relevant optimization techniques, applies improvements, evaluates performance, and iterates until optimal results are achieved.
The MARCO framework operates through three core components working in a coordinated iterative loop:
  • Web-Search Engine: We integrate the Tavily search tool [tavily2024] to retrieve relevant optimization techniques from scholarly sources, including acm.org, ieee.org, arxiv.org, and researchgate.net. The search engine is configured with adjustable parameters including search_depth (which controls the thoroughness of each search), include_answer (which determines whether to include direct answers or just source links), and the number of sources retrieved per query. This ensures that the optimizer agent has access to diverse and current references for each optimization iteration.
    Figure 3: Web-Search Integration Architecture. Schematic demonstrating how external scholarly sources (IEEE, ACM, arXiv, ResearchGate) feed current optimization strategies and techniques into MARCO’s iterative optimization loop, bridging the knowledge gap between LLM training cutoffs and current HPC research.
  • Performance Evaluator Agent: This agent benchmarks the generated code across multiple metrics, including execution time, memory usage, algorithmic complexity, and computational cost. It provides structured feedback to the optimizer agent, enabling the system to iteratively refine solutions based on quantitative performance data. Performance results are presented to users in a tabular interface, summarizing performance trends across iterations, correctness evaluations, and complexity assessments.
The three components operate in a closed feedback loop: the optimizer generates improved code based on web-sourced techniques, the evaluator measures performance improvements, and this feedback guides subsequent optimization iterations until convergence or user-defined stopping criteria are met.
V-C Evaluation of MARCO
We evaluated MARCO-optimized code against baseline LLM-generated solutions using the LeetCode 75 dataset [leetcode75] and an additional set of 10 randomly selected hard problems. Critically, we compared the performance of MARCO with and without the web-search component, benchmarking results against Claude-3.5 Sonnet and GPT-4o to isolate the contribution of dynamic knowledge retrieval.
Across all evaluated problems, MARCO with the Tavily web-search integration consistently outperformed both Claude-3.5 Sonnet and MARCO without web-search enhancement. This demonstrates the substantial value of incorporating real-time, externally sourced optimization strategies, especially in the rapidly evolving HPC landscape where new architectures, libraries, and techniques frequently emerge. The web-search component enables MARCO to access optimization techniques published after the base LLM’s training cutoff, providing a significant competitive advantage.
The iterative refinement process contributed significantly to performance gains. On harder LeetCode problems, subsequent iterations generally achieved better results than initial attempts, with the final iteration showing an average improvement of 30.9% over the base MARCO system without web search. This improvement aligns with our hypothesis that complex optimization problems benefit from multiple rounds of technique application and performance-based refinement.
Figure 4 illustrates the runtime performance differences between MARCO with and without the Tavily web-search tool across the evaluated problem set, clearly demonstrating the value of dynamic knowledge integration in automated code optimization.
Figure 4: Web-Search Component Performance Impact. Runtime comparison showing performance differences between MARCO with and without the Tavily web-search tool on the LeetCode 75 problem set. The results demonstrate consistent improvements when current optimization techniques are dynamically integrated into the optimization process.
VI Discussion
VI-A Key Findings and Analysis
Our experiments demonstrate that MARCO substantially enhances execution speed, memory efficiency, and algorithmic optimization compared to baseline LLM-generated code. The most significant finding is that performance improvements scale with problem complexity: while MARCO shows modest gains on simple problems (21.2% improvement on easy problems), it achieves substantial improvements on complex tasks (51.9% improvement on hard problems). This scaling behavior aligns with our hypothesis that multi-agent optimization provides the greatest value for computationally demanding scenarios where sophisticated optimization techniques offer substantial benefits.
The integration of the web-search component consistently yields superior optimizations across all problem categories, enabling the system to overcome the inherent knowledge limitations of static pretrained models. By dynamically retrieving cutting-edge techniques from recent research publications, MARCO bridges the critical gap between model training cutoffs and the rapidly evolving landscape of HPC optimization. Our results show a 30.9% improvement when web search is enabled, demonstrating the substantial value of real-time knowledge integration in automated code optimization.
The iterative refinement process proves particularly effective for complex algorithmic problems, where multiple optimization strategies can be layered and refined based on performance feedback. This multi-step approach mirrors human expert behavior, where initial solutions are progressively improved through systematic analysis and technique application.
VI-B System Limitations and Trade-offs
Despite its advantages, MARCO introduces several important trade-offs that must be considered for practical deployment. First, the multi-agent architecture incurs computational overhead through iterative communication between agents and the transmission of intermediate results. This results in slower code generation compared to traditional single-pass LLM workflows, with generation time increasing approximately linearly with the number of optimization iterations.
Second, while the web-search component enables access to recent research, the system’s performance remains constrained by the underlying LLM’s ability to correctly interpret and implement advanced HPC techniques retrieved from scholarly sources. Complex optimization strategies may be misapplied or inadequately integrated, particularly when they require deep domain expertise or specialized implementation knowledge.
Third, MARCO’s current implementation faces several technical constraints. The Performance Evaluator Agent lacks native support for key HPC libraries, including Numba, Torch, and CUDA, significantly limiting optimization potential for GPU-accelerated workloads. Additionally, the system is currently restricted to Python, which, while offering a rich ecosystem of optimization tools, excludes important HPC languages such as C++, Fortran, and specialized domain-specific languages.
Finally, cost considerations present practical limitations. While MARCO’s approach avoids expensive fine-tuning, the iterative nature of the optimization process increases API usage costs. Our analysis shows higher per-query costs for more sophisticated models like GPT-4o, despite occasionally inferior performance compared to Claude 3.5 Sonnet, suggesting the need for careful model selection and cost-performance optimization.
VI-C Unexpected Findings and Implications
Our evaluation revealed several counterintuitive results that provide important insights for multi-agent system design. Most notably, standalone LLMs occasionally outperformed MARCO on simple problems, particularly those in the ”easy” category of the LeetCode 75 dataset. This occurs because simple problems inherently require minimal optimization, and MARCO’s multi-agent orchestration introduces unnecessary computational overhead for tasks that are already well-solved by basic implementations.
This finding has important implications for system design: effective automated optimization systems must include problem complexity assessment to determine when sophisticated optimization is warranted. Simple problems may be better served by direct LLM generation, while complex problems benefit from MARCO’s iterative refinement approach.
The scaling relationship between problem complexity and optimization benefit suggests that MARCO’s architecture is well-suited for its intended domain of high-performance computing, where problems are typically computationally demanding and benefit from sophisticated optimization techniques. This validates our design hypothesis that multi-agent systems provide the greatest value for complex, domain-specific tasks rather than general-purpose code generation.
VI-D Broader Implications and Future Directions
These findings have significant implications for the development of AI-assisted programming tools in specialized domains. The success of dynamic knowledge integration through web search suggests that hybrid approaches combining pretrained models with real-time information retrieval may be more effective than purely static or purely fine-tuned approaches for rapidly evolving technical domains.
Future work should address the identified limitations through several key directions: expanding language support beyond Python to include C++ and Fortran, integrating support for GPU acceleration libraries, developing more sophisticated problem complexity assessment mechanisms, and optimizing the cost-performance trade-offs in multi-agent architectures.
Additionally, evaluation on real-world HPC applications beyond algorithmic benchmarks would provide valuable insights into MARCO’s practical effectiveness for scientific computing workloads. The development of domain-specific evaluation metrics for HPC code quality, including parallelization efficiency and hardware utilization, represents another important research direction.
The demonstrated effectiveness of MARCO’s approach suggests promising applications to other specialized programming domains where current LLMs show limitations, including systems programming, embedded software development, and domain-specific scientific computing applications.
VII Conclusions and Future Work
VII-A Summary of Contributions
In this work, we addressed the significant limitations of current LLMs in generating optimized HPC code by introducing MARCO, a novel multi-agent system designed to iteratively refine code for improved performance, memory efficiency, and scalability. Our comprehensive evaluation demonstrates that MARCO effectively leverages external knowledge sources and iterative optimization loops to substantially outperform standalone LLMs, achieving a 14.6% average runtime reduction on the LeetCode 75 problem set and a remarkable 51.9% improvement on complex problems compared to Claude 3.5 Sonnet.
The integration of real-time web search capabilities proves particularly valuable, yielding a 30.9% performance improvement over the base MARCO system and enabling access to cutting-edge optimization techniques beyond the knowledge cutoff of pretrained models. These results establish MARCO as an effective approach to automated HPC code optimization that bridges the gap between static LLM capabilities and the rapidly evolving landscape of high-performance computing.
Our key contributions include: (1) a specialized multi-agent architecture that separates code generation, evaluation, and testing responsibilities; (2) an adaptive feedback mechanism that enables iterative code refinement; (3) real-time knowledge integration through web search of scholarly sources; and (4) a cost-effective implementation that avoids expensive fine-tuning while delivering superior performance.
VII-B Future Research Directions
Building upon these promising results, several important directions emerge for future development and research. We organize these opportunities into three primary categories: system enhancements, evaluation expansion, and broader applications.
VII-B1 System Enhancements
The most immediate improvements focus on enhancing inter-agent communication and expanding system capabilities. Currently, the Performance Evaluator Agent provides feedback solely through final performance metrics. Implementing richer bidirectional communication—including intermediate diagnostic information, specific bottleneck identification, and targeted optimization suggestions—could significantly improve the refinement process efficiency and effectiveness.
Technical infrastructure improvements represent another crucial development area. Integrating Docker-based execution environments would support a wider range of external libraries (including Numba, Torch, and CUDA) and enable more robust, cross-platform testing capabilities. Additionally, extending language support beyond Python to include C++, Fortran, and domain-specific languages would significantly broaden MARCO’s applicability to real-world HPC applications.
The optimization workflow itself offers substantial improvement potential. Rather than generating complete code variants for each iteration, MARCO could maintain and incrementally edit persistent code artifacts, enabling finer-grained improvements and reducing redundant computations. Incorporating reinforcement learning techniques, as demonstrated in systems like AlphaCode [li2022competition], could enable MARCO to learn from iterative feedback and improve optimization strategies over time [sun2023reinforcement, sun2024llm].
VII-B2 Evaluation and Validation
Future evaluation efforts should expand beyond algorithmic benchmarks to include real-world HPC applications across diverse scientific domains. Developing domain-specific evaluation metrics that capture HPC-relevant qualities—such as parallelization efficiency, hardware utilization, and scalability characteristics—would provide more comprehensive assessment of optimization effectiveness.
Long-term studies examining MARCO’s performance consistency across different hardware architectures and workload types would validate the generalizability of our findings. Additionally, comparative studies with other automated optimization approaches, including compiler optimizations and specialized fine-tuned models, would better position MARCO within the broader optimization landscape.
VII-B3 Broader Applications and Impact
The success of MARCO’s multi-agent approach suggests promising applications to other specialized programming domains where current LLMs show limitations, including systems programming, embedded software development, and domain-specific scientific computing. The demonstrated effectiveness of dynamic knowledge integration through web search may prove valuable for rapidly evolving technical domains beyond HPC.
Furthermore, integrating persistent communication history across optimization sessions could reduce token usage costs and enable more sophisticated multi-session optimization pipelines, making the approach more practical for large-scale deployment in research and industrial settings.
VII-C Concluding Remarks
MARCO represents a significant advancement in automated HPC code optimization, demonstrating that multi-agent orchestration combined with real-time knowledge retrieval can substantially improve upon traditional single-model approaches. By effectively bridging the gap between static pretrained LLMs and the rapidly evolving landscape of high-performance computing, MARCO establishes a foundation for scalable, cost-efficient, and accessible optimization pipelines.
The demonstrated performance improvements, particularly on complex problems where sophisticated optimization provides the greatest benefit, suggest that MARCO’s approach addresses a genuine need in the HPC community. As computational demands continue to grow across scientific and industrial domains, automated optimization tools like MARCO have the potential to significantly lower barriers to high-performance programming and accelerate innovation in computationally intensive applications.
The path forward involves continued development of the technical capabilities outlined above, broader evaluation across real-world applications, and exploration of the approach’s applicability to other specialized programming domains. Through these efforts, MARCO’s multi-agent optimization paradigm may contribute to making high-performance computing more accessible and effective for the broader scientific and engineering communities.